[Hallc_running] [New Logentry] RC Report - 7/19/22

arunts at jlab.org arunts at jlab.org
Tue Jul 19 10:45:02 EDT 2022

Logentry Text:
We continue to take production data on LH2 at 65 uA (until we reach 32k pions) and dummy at 40 uA. We have a configuration change that's coming up at 11:00 am where we switch to Q^2 = 8.5 GeV^2 data point. Spectrometer rotation should be enabled already since Steve Lassiter left it enabled after the previous configuration change. For maintenance day tomorrow, we plan to address a number of tasks in the hall.
Plan for tomorrow:
-Beam stops at 6:30 am, RadCon will perform a full survey and the hall will go into controlled access.
-Walter + tech -->  pump on the HMS transfer line (few hours).
-Heidi --> verify SHMS N2 meter calibration settings, and work with Radcon to check levels in the SHMS Shield hut (few hours). 

Let's have a brief RC meeting today at 11:00 AM EST.

Meeting ID: 160 710 0885 - Passcode: 595071


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/4017460
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20220719/36685ee5/attachment.html>

More information about the Hallc_running mailing list